# Robust Design and Evaluation of

### v5

@@ 4.1 Bounds on positive class performance
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator, and define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\left(1+(1-D_{i})(\underline{\Gamma}-1)\right)\widehat{\delta}(X_{i})]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
@@
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}\right),$
The analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Proposition 4.1 establishes that the error of our estimator for the true bounds on positive class performance consists of two terms. The first term consists of the doubly-robust product of nuisance function errors. The second term depends on the root mean squared error of the estimator $\hat{\delta}(\cdot)$ for the nuisance function $\mu_{1}(\cdot)$ that enters the constraints. In other words, plugging in estimated constraints only affects the convergence rate of our estimator through the root mean squared error of the estimated constraints. Since this result applies to any choice of estimator $\hat{\delta}(\cdot)$ for the nuisance function, it can be combined with known results on the convergence rates of off-the-shelf nonparametric regression procedures. Of course, care must be taken in the choice of nonparametric regression procedure to enforce that the constraints $\widehat{\Delta}_{n}$ are feasible.
Proving this result involves two key steps. First, we establish that optimization over the set of possible confounding function satisfying Equation (6) is in fact equivalent to optimization over the class of monotone, non-decreasing functions. Such a reduction has been noted before in other sensitivity analysis models such as AronowLee(2013); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21). Second, we establish that optimization over the set of estimated bounding functions $\widehat{\Delta}_{n}$ introduces an error that is bounded above by the mean square estimation error in the bounding functions.
As one example of how Proposition 4.1 may be applied in conjunction with convergence rates on nonparametric regression procedures, we consider the case in which we estimate the nuisance function $\mu_{1}(\cdot)$ using our pseudo-outcome regression procedure. On the second fold, suppose we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding a particular estimator $\widehat{\delta}(\cdot)$. Our analysis in Section 3 implies that the use of pseudo-outcome regression based on efficient influence functions and sample splitting is sufficient to control bias from the estimation of nuisance functions that appear in the constraints, and it delivers the following result.
@@
Under the same conditions as Proposition 4.1, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), then
@@
4.2 Bounds on overall performance
We next estimate the bounds on overall performance. Under nonparametric outcome bounds, the upper bound on overall performance can be written as
The lower bound can be rewritten analogously. The bounds on overall performance are linear functionals of known functions of the data and identified nuisance parameters $\eta=\left(\mu_{1}(\cdot),\pi_{0}(\cdot)\right)$. We therefore construct debiased estimators for the bounds on overall performance using standard arguments based on efficient influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview). For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting, characterize their convergence rates and provide conditions under which they are jointly asymptotically normal.
To estimate the bounds on the overall performance, we randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\widehat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
for each observation $i$ in the $k$-th fold, where again $\phi_{\mu,i}(\eta)$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta)$ is the uncentered efficient influence function. These can be derived using standard arguments (e.g., Hines_2022). We then average over all observations
$\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$
where $K_{i}$ is the fold containing observation $i$. Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
Proposition 4.2.
Define $R_{1,n}^{k}:=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$. Then,
[…]
The rate condition required for the estimators to be jointly asymptotically normal will be satisfied if all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$, which is the familiar condition from the debiased machine learning literature. A wide-suite of nonparametric regression methods or modern machine learning based methods can therefore be used to estimate the nuisance functions.
In Online Appendix C.2, we construct a consistent estimator of the asymptotic covariance matrix in Proposition 4.2. The joint asymptotic normality of our estimators combined with the consistent estimator of the asymptotic covariance matrix imply that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods for inference under partial identification (e.g., ImbensManski(04); Stoye(09)--PICI).
5 Monte Carlo Simulations
We now illustrate that the theoretical guarantees for our proposed estimators provide guidance about their behavior in finite samples.
We simulate training and evaluation datasets satisfying Equation (6). In particular, we draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$. We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
for some $d_{\mu}\in\{1,\ldots,d\}$ and $10<\Gamma_{true}<$ 1 We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
5.1 Finite sample behavior of estimated conditional probability bounds
We first compare three estimators for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes based on the efficient influence function. Our estimator and the oracle procedure are constructed using a single split of the evaluation data. The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso. Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle procedure, whereas the plug-in estimator may inherit errors from the estimation of the nuisance functions.
Figure 1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
[…]
Notes: This figure plots the average integrated mean square error of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$ across Monte Carlo simulations. We report results for $n\in\{1000,2500,5000\}$ and $d\in\{100,500\}$. The results are computed over 1,000 simulations.
Across 1000 simulated datasets of varying size $n\in\{1000,2500,5000\}$ and dimension $d\in\{100,500\}$, Figure 1 reports the average integrated mean square error of our estimator $\widehat{\overline{\mu}}(\cdot)$, the oracle learner $\widehat{\overline{\mu}}_{oracle}(\cdot)$, and the plug-in learner for the true bound $\overline{\mu}^{*}(\cdot)$. We find analogous results when analyzing the lower bound. As $n$ grows large, the average integrated mean square error of our estimator converges towards zero alongside the integrated mean square error of the oracle learner as expected. While it is competitive for smaller sample sizes ($n=1000$), the average integrated mean square for the plug-in learner is constant across sample sizes $n$ and dimension of the features $d$. As $n$ grows larger, the plug-in learner performs poorly – for example, at $n=5000$ and $d=500$, our estimator’s integrated mean square error is only 1.9% larger than that of the oracle learner, whereas the plug-in learner’s integrated mean square error is 18.4% larger. In contrast, by leveraging both sample-splitting and pseudo-outcomes based on efficient influence functions, our estimator quickly converges to the true bound and improves upon simple plug-in estimation approaches.
5.2 Finite sample behavior of estimated predictive performance bounds
To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$. We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and nonparametric outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
$n$ Estimate SD Bias
500 0.545 0.016 0.001
1000 0.544 0.010 0.000
2500 0.544 0.007 0.000
(a) Upper bound on true positive rate
$n$ Estimate SD Bias
500 0.504 0.015 -0.003
1000 0.504 0.010 -0.004
2500 0.504 0.007 -0.003
(b) Lower bound on true positive rate
Table 1: Bias properties of estimators for the bounds on the true positive rate.
[…]
Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta),\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the bounds on the true positive rate and the standard deviation of our estimators across simulations. We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$. The results are computed over 1,000 simulations.
We explore how well our proposed estimators recover the bounds on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$. We construct our estimators of the bounds using three splits of the evaluation data. We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold. We trim the estimated bounding functions so that the constraints $\widehat{\Delta}_{n}$ are non-empty. Table 1b reports the average bias of our estimators for the bounds. Their bias quickly converges to zero as the size of the evaluation data grows large, illustrating our analysis of their convergence rates in Proposition 4.1. We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding. Figure 2 reports the simulation distribution of our proposed estimators as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$. Across alternative assumptions of unobserved confounding, our proposed estimators are approximately centered at the bounds on the true positive rate.
Figure 2: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
[…]
Notes: This figure illustrates box-plots (red) for the distribution of estimators of the bounds on the true positive rate as $\underline{\Gamma}=1/\tilde{\Gamma}$, $\overline{\Gamma}=\tilde{\Gamma}$ varies. The dashed black lines show the true upper and lower bounds for each value of $\tilde{\Gamma}$. Results are reported for $n=1000$ and computed over 1,000 simulations.
As mentioned in Section 4.2, our proposed estimators for the overall performance bounds follow a standard debiased construction. We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties. For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table 2b reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2. Our proposed estimators are approximately unbiased for the true bounds. Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$. Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage.
@@ 6 Empirical Application to Credit Risk Scores
Consider a financial institution using a credit risk score $s(\cdot)$ to predict whether an application will default based on application-level covariates. Our goals are to design a new credit risk score that accounts for possible unobserved confounding and to audit the predictive performance of a credit risk score. We observe data on past applications, but only observe whether a past application defaulted on the loan if the application was “funded” (i.e., approved by the financial institution and the offered terms were accepted by the applicant). The funding decision $D_{i}\in\{0,1\}$ may be subject to unobserved confounding as applicants may differ in unobserved ways that jointly affect their default risk and likelihood of accepting an offered loan. For example, an applicant’s decision to accept an offered loan may depend on whether they secured another credit offer at a competing financial institution.
We use data from Commonwealth Bank of Australia (“CommBank”), a large financial institution in Australia, on 372,346 applications submitted from July 2017 to July 2019. As discussed in CostonEtAl(21)-Rashomon, personal loans are paid back in monthly installments, and used, for example, to purchase a used car, refinance existing debt, pay for funeral and wedding expenses among many other purposes. In the time period considered, personal loan sizes included amounts up to AU\$50,000 with a median of AU\$10,000, and the median offered interest rate was 14.9% per annum in our sample of applications. We observe rich application-level covariates such as the applicant’s reported income, the applicant’s occupation, and their credit history at CommBank. We only observe whether an applicant defaulted on the personal loan within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$. In our sample of applications, approximately one-third of the applications were funded, and 2.0% of all funded loans defaulted within 5 months.

### v6

@@ 4.2.1 Estimation via Linear Fractional Programming:
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists $M<\infty$ such that $\|\beta_{0}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (v) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
@@
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right).$
Proposition 4.2 reveals the structure of how estimation errors affect our estimator. The first term $O_{\mathbb{P}}(1/\sqrt{n})$ reflects sampling error that would arise even if the nuisance functions $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ were known. The second term $O_{\mathbb{P}}(R_{1,n})$ reflects errors from estimating the nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ that appear in the numerator and denominator of the linear fractional program. These enter through a doubly robust product, and this behavior is inherited from the influence function structure underlying the numerator and denominator. The final term captures errors from estimating the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ that define the constraint set $\Delta$. This error enters through the root mean square error. This error can be controlled using standard nonparametric regression techniques, and researchers can trade off bias and variance in familiar ways.
For this reason, Proposition 4.2 establishes a “partial” double robustness: given estimated constraints, our estimator is doubly robust to errors in the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$. Constraint estimation errors enter additively through RMSE. Achieving full double robustness — where constraint estimation errors would also enter with product structure — remains an open problem, which we discuss in Remark 4. Remark 5 describes a simple alternative approach that obtains fully doubly-robust estimators for non-sharp bounds on positive class performance.
Proposition 4.2 and our estimator are practically valuable. Researchers can use state-of-the-art machine learning methods for estimating the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$, while separately controlling the constraint estimation error through their choice of estimator for the bounding functions. The error decomposition applies to any estimator of the bounding functions with favorable $L_{2}(\mathbb{P})$ properties, and we discuss alternative strategies and their tradeoffs in Section 4.2.2. This modularity is a key practical advantage of our framework.
Remark 3 (Connection to existing work).
Our analysis extends existing work on linear fractional programs, including AronowLee(2013), MiratrixWagerZubizaretta(18), KallusZhou(18)-confoundingrobust, and ZhaoSmallBhattacharya(19). These papers analyze linear fractional programs that arise when bounding population means or treatment effects under alternative sensitivity analysis models. Our analysis advances in two key respects: (i) we establish partial double robustness for objective nuisances – a property not achieved in prior work; and (ii) we provide a unified framework that accommodates multiple bounding strategies (observed outcome bounds, proxy outcome bounds, instrumental variable bounds) through a modular approach to constraint estimation.
Remark 4 (Paths toward full double robustness).
The proof of Proposition 4.2 reveals that the optimizer in the linear fractional program can be characterized as having a threshold structure in $\beta_{0,i}$. The twist is that the cutoff determining the threshold is itself not fixed but rather a functional of the estimated constraints. Consequently, estimation error in the constraints affects not only which functions are feasible but also where the optimal threshold lies. While related, recent work on conditional linear programs (benmichael2025partialidentificationconditionallinear) and aggregated intersection bounds (semenova2023adaptive) are not directly applicable. While the linear fractional program can be reduced to a linear program through the Charnes-Cooper transformation, the transformation uses a global normalization that couples all observations together, whereas existing work exploits conditional separability across $X$-cells. We leave full resolution of this question for future work and view our partial double robustness result as an important first step.
Remark 5 (Estimation of non-sharp bounds).
Our estimator targets sharp bounds on positive class performance, but an alternative approach that restores double robustness at the cost of sharpness is possible. The idea is straightforward: separately estimate the numerator and the denominator using overall performance estimators (Section 4.1) and then combine via the delta method. Concretely, define
$\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)=\frac{\sup_{\delta\in\Delta}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
Clearly, $\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)\geq\overline{\mbox{perf}}_{+}(s;\beta)$. The numerator and the denominator can be separately estimated using the methods in Section 4.1. This delivers: (i) full double robustness with respect to all nuisances, and (ii) standard inference with straightforward variance estimation. The cost is that we target non-sharp bounds. Whether this trade-off is attractive depends on the application.
Remark 6 (Interpretation of non-empty constraint sets).
Proposition 4.2 assumes the estimated constraint set $\widehat{\Delta}_{n}$ is non-empty. If finite-sample variability leads to $\widehat{\underline{\delta}}(x)>\widehat{\overline{\delta}}(x)$ for some $x\in\mathcal{X}$, researchers can either: (i) clip the bounds to enforce feasibility, which might introduce bias captured by the root mean square terms in Proposition 4.2; or (ii) restrict attention to the feasible covariate region $\tilde{X}_{n}=\{x:\widehat{\underline{\delta}}(x)\leq\widehat{\overline{\delta}}(x)\}$, thus redefining the estimand. The latter approach is analogous to trimming observations with extreme estimated propensity scores in causal inference, though this may complicate analysis (e.g., HotzEtAl(09); Sasaki_Ura_2022).
4.2.2 Estimation of the Bounding Functions
The error decomposition in Proposition 4.2 applies to any estimator of the bounding functions with favorable mean square error properties. This modularity gives researchers flexibility. We discuss two natural strategies and their relative merits, focusing on observed outcome bounds for concreteness.
The simplest approach directly substitutes nuisance estimates into the bounding function formulas. For observed outcome bounds, this means $\widehat{\underline{\delta}}(x)=(\underline{\Gamma}-1)\widehat{\mu}_{1}(x)$ and $\overline{\delta}(x)=(\overline{\Gamma}-1)\widehat{\mu}_{1}(x)$. The mean square error properties of the estimated bounding functions are then inherited from those of $\widehat{\mu}_{1}(x)$. This approach requires no additional estimation beyond the first-stage nuisances. An alternative approach is to construct influence-function-based pseudo-outcomes and regress them on covariates using the full sample (i.e., our pseudo-outcome regression procedure in Section 3). We develop this approach below as it may offer bias-variance advantages in practice. Ultimately, both strategies are valid under our framework.
To understand why pseudo-outcome regression may be advantageous, consider that $\widehat{\mu}_{1}(x)$ is typically constructed using only observations with $D_{i}=1$. Consequently, standard regression procedures target the mean square error conditional on $D_{i}=1$ However, the error bound in Proposition 4.2 depends on the unconditional mean square error. These two objectives could lead to different bias-variance tradeoffs when the selected population ($D_{i}=1$) is small relative to the full population or there is covariate shift between the selected and unselected populations. In such settings, an estimator that directly targets the unconditional MSE may achieve better performance. Pseudo-outcome regression accomplishes this by constructing unbiased pseudo-outcomes using observations from both the selected and unselected populations, then regressing these pseudo-outcomes on the full sample.
Estimation Procedure for Observed Outcome Bounds:
To illustrate how Proposition 4.2 may be applied, we estimate bounding functions using our pseudo-outcome regression procedure. This illustrates how our partial double robustness result (Proposition 4.2) can be combined with analyses of nonparametric regression estimators — in particular, leveraging our oracle inequality for pseudo-outcome regression (Appendix A).
[…]
@@ 4.2.2 Estimation of the Bounding Functions
Under the same conditions as Proposition 4.2, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), then
@@ 5 Empirical Application to Credit Risk Scores
We validate the finite-sample performance of our estimators through Monte Carlo simulations in Appendix F. We now apply our framework to credit risk prediction, where financial institutions use scores $s(\cdot)$ to predict default risk. We observe data on past loan applications but only observe defaults $Y_{i}=D_{i}Y_{i}^{*}$ for funded applications ($D_{i}=1$). It is implausible to assume unconfounded funding decisions. Applicants who accept loan offers may differ systematically from those who decline in ways that affect their default risk. [^7] Lending institutions face two distinct inference problems. First, ”policy rejects” with $\mathbb{P}(D_{i}=1\mid X_{i})=0$ (i.e., overlap violations) due to underwriting rules require extrapolation-based methods that are beyond our scope. Second, even among fundable applicants with $\mathbb{P}(D_{i}=1\mid X_{i})>0$, those accepting versus declining offers may differ in unobserved ways — the selective labels problem we address. Extending our framework to handle policy rejects is an important direction for future research.
We analyze 372,346 personal loan applications submitted to Commonwealth Bank of Australia from July 2017 to July 2019. These loans are repaid in monthly installments and used to purchase vehicles, refinance debt, or cover major expenses (CostonEtAl(21)-Rashomon). Loan amounts ranged up to AU\$50,000 (median AU\$10,000) with a median offered interest rate of 14.9% per annum. We observe rich application-level covariates including reported income, occupation, and credit history at CommBank. We only observe whether an applicant defaulted within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$. Approximately one-third of applications were funded, with a 2.0% default rate among funded loans.
We consider two exercises. First, we construct credit risk scores robust to varying confounding assumptions. The resulting algorithms produce substantially different risk rankings that could affect credit access. Second, following recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate how a benchmark score’s accuracy varies across income groups. While an analysis ignoring confounding suggests better performance for higher-income applicants, this disparity disappears under plausible confounding levels.